OpenStudy (anonymous):

Find the average rate of change of f: (a) from 1 to 2 f(x)=8x^2-x

OpenStudy (anonymous):

i got an answer of 30 but the correct answer says it's 23

OpenStudy (anonymous):

@jim_thompson5910 could you check this for me please?

jimthompson5910 (jim_thompson5910):

what are f(2) and f(1) ?

OpenStudy (anonymous):

f(2)=30 and f(1)=0

jimthompson5910 (jim_thompson5910):

f(2) = 30 is correct f(1) = 0 is not correct

OpenStudy (anonymous):

oh ok i know what i did wrong

OpenStudy (anonymous):

are you available to help me with one more problem?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Sketch the graph of each function. a.) f(x)=l x l b.) f(x)=\[\sqrt[3]{x}\] c.)f(x)=\[\sqrt{x}\] d.)f(x)=1/x

OpenStudy (anonymous):

**Be sure to label atleast three points

jimthompson5910 (jim_thompson5910):

do you know how to generate a point for any function?

OpenStudy (anonymous):

i don't think i do

jimthompson5910 (jim_thompson5910):

say x = 0 f(x) = |x| f(0) = |0| ... replace every x with 0 f(0) = 0 ... take the absolute value of 0 to get 0 so one ordered pair is (0,0) since x = 0 and y = 0

jimthompson5910 (jim_thompson5910):

what is another ordered pair on y = |x| ?

OpenStudy (anonymous):

(2,2)?

jimthompson5910 (jim_thompson5910):

good, what are 2 more (use one positive x value and a negative x value)

OpenStudy (anonymous):

(-3,3) and (-1,1)

jimthompson5910 (jim_thompson5910):

so if you generate enough points, you'll get an idea of what the graph looks like

jimthompson5910 (jim_thompson5910):

connecting all of the points will generate the line or curve of the function

jimthompson5910 (jim_thompson5910):

the more points you generate, the more accurate the graph

OpenStudy (anonymous):

okay and that goes for all 4 functions correct?

jimthompson5910 (jim_thompson5910):

yes, any function for that matter

OpenStudy (anonymous):

okay, Thankyou!

jimthompson5910 (jim_thompson5910):

you're welcome